# Effective action for $\phi^{4}$-Yukawa theory via 2PI formalism in the inflationary de Sitter spacetime

### v2

@@ 2 The basic set up
where $a(t)=e^{Ht}$ or $a(\eta)=-1/H\eta$ is the de Sitter scale factor and $H=\sqrt{\Lambda/3}$ is the de Sitter Hubble rate, with $\Lambda$ being the cosmological constant. We have the range $0\leq t<\infty$, so that $-H^{-1}\leq\eta<0^{-}$. Note that any temporal level of the initial hypersurface can be achieved as we wish, by exploiting the time translation symmetry (along with a scaling of the spatial coordinates) of de Sitter.
@@
where ∇̸$\not{\nabla}=\tilde{\gamma}^{\mu}\nabla_{\mu}∇$, and $\tilde{\gamma}^{\mu}$ and $\nabla_{\mu}$ are respectively the curved space gamma matrices and the spin covariant derivative. Also, since we are working with the mostly positive signature of the metric, we will take the anti-commutation relation for the $\gamma$-matrices as
$[\tilde{\gamma}_{\mu},\tilde{\gamma}_{\nu}]_{+}=-2g_{\mu\nu}\,{\bf I}_{d\times d}=-2a^{2}\eta_{\mu\nu}\,{\bf I}_{d\times d}$ (3)
@@
Let us now come to the issue of the scalar field. The exact propagator for a massive scalar can be seen in 98 of B. Although we will chiefly work with a massive scalar in this paper (note that a background field makes the scalar effectively massive, even if the scalar has zero rest mass), we also wish to additionally compute some vacuum graphs for a massless and minimally coupled scalar field (A), which is of special interest in de Sitter. The propagator for a massless and minimally coupled scalar in desitter spacetime reads [17],
@@ 3.1 Warm up – a scalar with $\phi^{4}$ self-interaction
The term appearing in the last term on the right hand side of 17, $2\delta i\Gamma_{2}/\delta iG(x,x^{\prime})$ is basically $-i$ times the ${\cal O}(\lambda)$ 1PI self energy. This whole last term explicitly reads
@@
Let us now come to the non-perturbative renormalisation scheme. From 104 (B) valid for an arbitrary massive scalar, we write for the sake of brevity
@@
We next plug in 20 into 23. For a consistent renormalisation, we first infer that
@@
[…]
@@
$\delta\lambda_{1}=3\delta\lambda_{2},\qquad\delta m_{2}^{2}=\delta m_{1}^{2},\qquad\delta\xi_{1}=0$ (31)
@@
[…]
@@
[…]
@@
Figure 2: Variation of the effective potential 35, with respect to the background field $v$, for negative rest mass squared $m_{0}^{2}\sim-0.063H^{2}$. The bar over the quantities denote scaling with respect to appropriate power of the de Sitter Hubble rate, $H$. The blue, red and black curves respectively correspond to $\lambda$ values $0.10$, $0.11$ and $0.15$. This non-trivial feature was reported first in [76, 77, 78].
@@ 3.2 Inclusion of the Yukawa interaction
Thus $iS_{l}(x,x^{\prime})$ is simply the propagator for a fermion with mass $M_{\rm dyn,eff}$. Now, for the local part of the self energy, we will need the fermion propagator only at small scales, so that we obtain a $\delta$-function in the integrand in the self energy integral of 42, shrinking the loop to a single point, and giving rise to the dynamically generated fermion rest mass. We take the fermion to be light. Then the part of the propagator relevant for our present purpose can be red off from 108 of B ([66]),
@@
[…]
Using 44, let us now compute the self energy term appearing in 42. Using the expression 101 for the scalar propagator and 83, we have
@@
Note that $iS_{l}(x,x^{\prime})$, 44, also yields non-local terms via the logarithms of 83, which we are not considering here. We now plug the above expression into 42. The divergence associated with the first term on the right hand side of 45 can be tackled by a fermion field strength renormalisation counterterm, $\delta Z_{f}$. Following [61], we write
@@
[…]
@@
Thus for example for the range $g\sim 0.1-0.5$, we see that the logarithm term contributes at most $0.7-17\%$ to $M_{0}+gv$. Due this reason we shall take $M_{\rm dyn,eff}\simeq M_{0}+gv$ in the following, and will ignore the dynamically generated fermion mass.
@@
where we have used 20, 21, 111. We will also take $M_{\rm dyn,eff}\simeq M_{0}+gv$, as argued above. We also have suppressed the scalar field strength renormalisation counterterm as it will not be necessary for our present purpose.
@@
It is clear that renormalising the Yukawa contribution entirely or keeping it entirely corresponds respectively to $k=0$ or $k=1$ in the above equation. The above decomposition corresponds to the dynamical effective scalar mass squared
@@
Note that $\delta\lambda_{2}\neq\delta\lambda_{3}$ here, unlike the earlier case of pure scalar field theory. The structure of the counterterms appearing above suggest various non-trivial self energy and vertex corrections and their resummation. For example, the last term on the right hand side in the expression for $\delta\lambda_{2}$ corresponds to the renormalisation of the box diagram corresponding to the scalar four point function in the Yukawa theory. On the other hand, the second term on the right hand side of the same corresponds to the resummed renormalisation of the quartic scalar vertex correction due to the insertion of the Yukawa vertices, 6.
@@
Using next 20, 59 and 111 of C, we obtain for 62 after a little algebra
@@
Now we note in the above expression that the most problematic divergence appears in the penultimate line, given by
@@
Given the structure of $f_{\rm fin}$, 20, the above divergence cannot be absorbed by any counterterm we know of. This is an overlapping divergence and is absent in the traditional computations. Recall also that we do not have the freedom here to add any further graph in the effective action to attempt a cancellation. Putting things together, we make the choice $k=-1$ in 58, 60 and 64, so that the above overlapping divergence gets cancelled by the last term of the fourth line of 64. It is the necessity of cancellation of this divergence that led us to introduce the constant $k$. Any choice other than $k=-1$ (like $k=1$ or $k=0$) would not have served our purpose. Note also that this problem would not have arisen if the sign in front of the last term on the right hand side of the effective action 62 was opposite. However, we have checked that there is no error in any sign.
@@
We next compute for 62 using the unrenormalised coincidence limit expressions 20 and 108,
[…]
where we have abbreviated from 108,
@@
$+\int a^{d}d^{d}x\left[\left\{\frac{\lambda((m_{0}^{2}f_{d}+H^{2}f^{\prime}_{d})-C(M_{0}^{2}+H^{2})f_{d})^{2}}{2^{3}(1+\lambda f_{d}/2)}-\frac{m_{0}^{2}}{2}\left(m_{0}^{2}f_{d}+2H^{2}f^{\prime}_{d}\right)+\frac{M_{0}^{2}H^{-\epsilon}}{2^{1-\epsilon/2}\pi^{2-\epsilon/2}}\left(\frac{M_{0}^{2}}{2}\left(-\frac{2}{\epsilon}-\frac{3}{2}+\gamma\right)\right.\right.\right.$
@@
where $M_{\rm dyn,eff}$ and $F_{\rm fin}$ is given by 67, $m^{2}_{\rm dyn,eff}$ by 59 (with $k=-1$), $f_{\rm fin}$, $f_{d}$ and $f^{\prime}_{d}$ by 21 and 22 and $C$ is defined in 63. The first two lines of the above equation are finite. The fifth line gives the divergences corresponding to the quartic self interaction for mass and coupling constant renormalisation. The remaining lines give the divergences corresponding to the Yukawa or overlapping Yukawa-$\phi^{4}$ interactions, and they all vanish in the absence of the Yukawa coupling. When 68 is plugged into 62, the terms proportional to $v$, $v^{2}$, $v^{3}$ and $v^{4}$ respectively give the tadpole ($\gamma_{\rm tad}$), the mass renormalisation ($\delta m_{1}^{2}$), the cubic coupling ($\delta\beta_{1}$) and the quartic coupling ($\delta\lambda_{1}$) counterterms. We set $\delta\xi_{1}=0$. Note also that for vanishing fermion rest mass, we have $\delta\beta_{1}=0=\gamma_{\rm tad}$. Finally, the third and fourth lines are field independent and hence must be absorbed in a cosmological constant counterterm in the gravitational action.
@@
where we have absorbed further some field independent constant terms into redefinition of the cosmological constant in the gravitational action. 70 is the final result of this paper. Note that owing to our estimation of the secular logarithm, the above effective action is valid only towards the end of the inflation. At some earlier stages, one may just ignore the $\ln a$ term.
@@
[…]
@@
we have from 67, 59 (with $k=-1$) the leading expression
@@
Although this high field limit 72 is similar to that of the standard one loop effective action, the coefficients of the $v^{4}\ln v$ term contains contribution beyond one loop. Note from 73 that if we want $\lambda$ to be less than unity, we must have $g\lesssim 0.3$. This is consistent with our argument $M_{\rm dyn,eff}\simeq M_{0}+gv$, we made earlier for the fermion (cf, the discussion below 54). Finally, we also note that $|v|/H\gg 1$ can also be interpreted as the flat spacetime $(H\to 0)$ limit of 71. The constraint 73 ensures that 72 has the standard Coleman-Weinberg spontaneous symmetry breaking feature, as depicted in 7.
@@
We have depicted the behaviour of the non-perturbative effective potential for massless fields, 71, in 8, with $\lambda>16g^{2}$, 73. Let us compare them with the pure scalar field theory results, 2, 3. For a scalar with zero or positive rest mass squared, 3 are rather similar to the tree level potential. 8 on the other hand, shows behaviour qualitatively similar to that of 2, which corresponds to negative rest mass square for the scalar field in a pure $\phi^{4}$-theory. 9 shows the feature of the non-perturbative effective potential with $\lambda=16g^{2}$. Note that for this equality, the $H\to 0$ limit of the effective potential is just the tree level quartic potential, which is very different from that of 9. The above discussion shows that the non-perturbative quantum effects can bring in novel features in spacetimes like the de Sitter. We also recall that for very high energy phenomenon occurring at very small length scales one usually ignores gravity and spacetime curvature. Then the comparison of our non-perturbative results with that of the $H\to 0$ limit also gives us one example of how quantum physics can be very different at short and large scales.

### v3

@@ 2 The basic set up
where $a(t)=e^{Ht}$ or $a(\eta)=-1/H\eta$ is the de Sitter scale factor and $H=\sqrt{\Lambda/3}$ is the de Sitter Hubble rate, with $\Lambda$ being the cosmological constant. We have the range $0\leq t<\infty$, so that $H^{-1}\leq\eta<0^{-}$. Note that any temporal level of the initial hypersurface can be achieved as we wish, by exploiting the time translation symmetry (along with a scaling of the spatial coordinates) of de Sitter.
@@
where $\not{\nabla}=\tilde{\gamma}^{\mu}\nabla_{\mu}$, and $\tilde{\gamma}^{\mu}$ and $\nabla_{\mu}$ are respectively the curved space gamma matrices and the spin covariant derivative. Also, since we are working with the mostly positive signature of the metric, we will take the anti-commutation relation for the $\gamma$-matrices as
$\tilde{\gamma}_{\mu},\tilde{\gamma}_{\nu}]_{+}=-2g_{\mu\nu}\,{\bf I}_{d\times d}=-2a^{2}\eta_{\mu\nu}\,{\bf I}_{d\times d}$ (3)
@@
Let us now come to the issue of the scalar field. The exact propagator for a massive scalar can be seen in 100 of B. Although we will chiefly work with a massive scalar in this paper (note that a background field makes the scalar effectively massive, even if the scalar has zero rest mass), we also wish to additionally compute some vacuum graphs for a massless and minimally coupled scalar field (A), which is of special interest in de Sitter. The relevant propagator such a scalar reads [17],
@@ 3.1 Warm up – a scalar with $\phi^{4}$ self-interaction
The term appearing in the last term on the right hand side of 17, $2\delta i\Gamma_{2}/\delta iG(x,x^{\prime})$ is basically i-i times the ${\cal O}(\lambda)$ 1PI self energy. This whole last term explicitly reads
@@
Let us now come to the non-perturbative renormalisation scheme. From 106 (B) valid for an arbitrary massive scalar, we write for the sake of brevity
@@
We next plug in 20 into 23. We first infer that
@@
$+$ $f_{\rm fin}^{2}.\frac{1}{8}\lambda^{2}f_{d}+H^{2}\left(m_{0}^{2}+\frac{1}{2}\lambda v^{2}+\frac{1}{2}\lambda f_{\rm fin}\right)f_{d}^{\prime}+\int dm^{2}_{\rm dyn,eff}f_{\rm fin}$
@@
$\delta\lambda_{1}=3\delta\lambda_{2},\qquad\delta m_{2}^{2}=\delta m_{1}^{2},\qquad\delta\xi_{1}^{2}=0$ (31)
@@
$\Gamma_{\rm 1PI}[v]_{\rm Ren.}=-\int a^{d}d^{d}x\left[\frac{1}{2}(\nabla_{\mu}v)(\nabla^{\mu}v)+\frac{1}{2}m_{0}^{2}v^{2}+\frac{\lambda}{4!}v^{4}-\frac{\lambda}{8}f_{\rm fin}^{2}+\frac{1}{2}\int d{m}^{2}_{\rm dyn,eff}f_{\rm fin}\right]$ (34)
@@
$\bar{V}_{\rm eff}(\bar{v})=\frac{V_{\rm eff}(v)}{H^{4}}=\left(\frac{1}{2}\bar{m}_{0}^{2}\bar{v}^{2}+\frac{\lambda}{4!}\bar{v}^{4}-\frac{\lambda}{8}\bar{f}_{\rm fin}^{2}-\frac{1}{2}\int d\bar{m}^{2}_{\rm dyn,eff}\bar{f}_{\rm fin}\right),$ (35)
@@
Figure 2: Variation of the effective potential 35, with respect to the background field $v$, for negative rest mass squared $m_{0}^{2}\sim-0.063H^{2}$. The bar over the quantities denote scaling with respect to appropriate power of the de Sitter Hubble rate, $H$. The blue, red and black curves respectively correspond to $\lambda$ values $0.10$, $0.11$ and $0.15$ respectively. This non-trivial feature was reported first in [76, 77, 78].
@@ 3.2 Inclusion of the Yukawa interaction
Thus $iS_{l}(x,x^{\prime})$ is simply the propagator for a fermion with mass $M_{\rm dyn,eff}$. Now, for the local part of the self energy, we will need the fermion propagator only at small scales, so that we obtain a $\delta$-function in the integrand in the self energy integral of 42, shrinking the loop to a single point, and giving rise to the dynamically generated fermion rest mass. We take the fermion to be light. Then the part of the propagator relevant for our present purpose can be red off from 110 of B ([66]),
@@
where $\not{\Delta x}$ contains contraction with respect to the flat space gamma matrices. Note also that since we are working in local approximation, the propagators appearing above are all Feynman propagators, and we do not need to consider the Wightman functions necessary for the in-in formalism described in A, just like the $\phi^{4}$-theory in the two loop Hartree approximation discussed earlier.
Using 44, let us now compute the self energy term appearing in 42. Using the expression 103 for the scalar propagator and 85, we have
@@
Note that $iS_{l}(x,x^{\prime})$, 44, also yields non-local terms via the logarithms of 85, which we are not considering here. We now plug the above expression into 42. The divergence associated with the first term on the right hand side of 45 can be tackled by a fermion field strength renormalisation counterterm, $\delta Z_{f}$. Following [61], we write
@@
where in $\not{\partial}$, contraction with respect to the flat space gamma matrix has been used. This immediately yields
@@
Thus for example for the range $g\sim 0.1-0.5$, we see that the logarithm term contributes at most $0.7-17\%$ to $M_{0}+gv$. Due this reason we shall take $M_{\rm dyn,eff}\simeq M_{0}+gv$ in the following, and will ignore the dynamically generated fermion mass. In other words, the following discussions will be valid only for small value of the Yukawa coupling. A formal resummed result for $M_{\rm dyn,eff}$ with generic $g$-values remains elusive to us so far.
@@
where we have used 20, 21, 113. We will also take $M_{\rm dyn,eff}\simeq M_{0}+gv$, as argued above. We also have suppressed the scalar field strength renormalisation counterterm as it will not be necessary for our present purpose.
@@
It is clear from the above expressions that either renormalising away the Yukawa contribution entirely or keeping it entirely corresponds respectively to $k=0$ or $k=1$ in the above equation. This would have been the simplest choice. However as we have mentioned above, doing so would lead to problem in cancelling divergence in the effective action, and as we will derive below, the unique choice is k=1k=-1. However until we prove it, let us keep $k$ generically. Note that k=1k=-1 partly takes away the Yukawa contribution, and the rest remains as finite contribution. As a physical consequence of the above renormalisation, the dynamical effective scalar mass squared acquires an Yukawa contribution given by,
@@
Recall that by the virtue of 53, $(g+\delta g_{3})^{2}/\epsilon$ is a finite quantity. Note also the appearance of the Fermion mass squared, $(M_{0}+gv)^{2}$ in the above expression.
@@
Note that $\delta\lambda_{2}\neq\delta\lambda_{3}$ here, unlike the earlier case of pure scalar field theory. The structure of the counterterms appearing above suggest various non-trivial self energy and vertex corrections and their resummation. For example, the last term on the right hand side in the expression for $\delta\lambda_{2}$ corresponds to the renormalisation of the box diagram corresponding to the scalar four point function in the Yukawa theory. On the other hand, the second term on the right hand side of the same corresponds to the resummed renormalisation of the quartic scalar vertex correction due to the insertion of the Yukawa vertices, 6. In particular, note from 60 that the value of $k$ can decide the diagrams generated. For example for $\delta\lambda_{2}$, $k=1$ omits the Yukawa box diagram totally, whereas $k=0$ omits the diagrams of 6. k=1k=-1 on the other hand, keeps all the diagrams corresponding to all the counterterms.
@@
Using next 20, 59 and 113 of C, we obtain for 62 after a little algebra
@@
In the above expression the most problematic divergence appears in the penultimate line, given by
@@
Given the structure of $f_{\rm fin}$, 20, the above divergence cannot be absorbed by any standard counterterm we know of. This is an overlapping divergence and is absent in the traditional computations. Recall also that we do not have the freedom here to add any further graph in the effective action to attempt a cancellation. Putting things together, we now make the choice
k=1,k=-1,
in 58, 60 and 64, so that the above overlapping divergence gets cancelled by the last term of the fourth line of 64. It is the necessity of cancellation of this divergence that led us to introduce the constant $k$ earlier. Any choice other than k=1k=-1 (like $k=1$ or $k=0$) would not have served our purpose. Note also that this problem would not have arisen if the sign in front of the last term on the right hand side of the effective action 62 was opposite. However, we have checked that there is no error in any sign. The most important consequence of this renormalisation is the particular form of the scalar dynamical mass, 59, which will contribute to the non-trivial characteristics of the effective potential as follows.
@@
We next compute for 62 using the unrenormalised coincidence limit expressions 20 and 110,
$\frac{1}{2}\int d^{d}xa^{d}\int dm^{2}_{\rm dyn,eff}\ iG(x,x)-\int d^{d}xa^{d}\int dM_{\rm dyn,eff}\ iS(x,x)$
$=\int d^{d}xa^{d}\left[\frac{m^{4}_{\rm dyn,eff}}{4}f_{d}+\frac{H^{2}m^{2}_{\rm dyn,eff}}{2}f^{\prime}_{d}+\frac{1}{2}\int dm^{2}_{\rm dyn,eff}f_{\rm fin}\right]$
$-\int d^{d}xa^{d}\left[\left(-\frac{2}{\epsilon}+\frac{3}{2}+\gamma\right)\frac{M^{2}_{\rm dyn,eff}H^{2-\epsilon}}{2^{1-\epsilon/2}\pi^{2-\epsilon/2}}+\left(-\frac{2}{\epsilon}-\frac{3}{2}+\gamma\right)\frac{M^{4}_{\rm dyn,eff}H^{-\epsilon}}{2^{2-\epsilon/2}\pi^{2-\epsilon/2}}+\int dM_{\rm dyn,eff}F_{\rm fin}\right],$ (66)
where we have abbreviated from 110,
@@
[…]
@@
where $M_{\rm dyn,eff}$ and $F_{\rm fin}$ is given by 67, $m^{2}_{\rm dyn,eff}$ by 59 (with k=1k=-1), $f_{\rm fin}$, $f_{d}$ and $f^{\prime}_{d}$ by 21 and 22 and $C$ is defined in 63. The first two lines of the above equation are finite. The fifth line gives the divergences corresponding to the quartic self interaction for mass and coupling constant renormalisation. The remaining lines give the divergences corresponding to the Yukawa or overlapping Yukawa-$\phi^{4}$ interactions, and they all vanish in the absence of the Yukawa coupling. When 68 is plugged into 62, the terms proportional to $v$, $v^{2}$, $v^{3}$ and $v^{4}$ respectively give the tadpole ($\gamma_{\rm tad}$), the mass renormalisation ($\delta m_{1}^{2}$), the cubic coupling ($\delta\beta_{1}$) and the quartic coupling ($\delta\lambda_{1}$) counterterms. We set $\delta\xi_{1}=0$. Note also that for vanishing fermion rest mass, we have $\delta\beta_{1}=0=\gamma_{\rm tad}$. Finally, the third and fourth lines are field independent and hence must be absorbed in a cosmological constant counterterm in the gravitational action.
@@
where we have absorbed further some field independent constant terms into redefinition of the cosmological constant in the gravitational action. 70 is the final result of this paper. Note that owing to our estimation of the secular logarithm, the above effective action is valid only towards the end of the inflation. At some earlier stages, one may just ignore the $\ln a$ term. We will briefly come back to the issue of $\ln a$ towards the end of this Section, and will argue that use of a recently proposed renormalisation group inspired autonomous equation to assign a non-perturbative value to this secular logarithm yields qualitatively similar result.
@@
$V_{\rm eff,loc}=$ $\frac{\lambda v^{4}}{4!}-\int dM_{\rm dyn,eff}F_{\rm fin}+\frac{1}{2}\int dm^{2}_{\rm dyn,eff}f_{\rm fin}-\frac{\lambda f^{2}_{\rm fin}}{8}+\frac{0.7C}{4}\left(\lambda-2Cg^{2}\right)v^{2}\left(H^{2}+g^{2}v^{2}\right)$ (71)
@@
we have from 67, 59 (with k=1k=-1) the leading expression
@@
Although this high field limit 72 is similar to that of the standard one loop effective action, the coefficients of the $v^{4}\ln v$ term contains contribution beyond one loop. Note from 73 that if we want $\lambda$ to be less than unity, we must have $g\lesssim 0.3$. This is consistent with our argument $M_{\rm dyn,eff}\simeq M_{0}+gv$, we made earlier for the fermion (cf, the discussion below 54). Finally, we also note that $|v|/H\gg 1$ can also be interpreted as the flat spacetime $(H\to 0)$ limit of 71. The constraint 73 ensures that 72 has the standard Coleman-Weinberg spontaneous symmetry breaking feature, as depicted in 7. Also note that since the violation of the upper bound of 73 makes $V_{\rm eff,loc}$ unbounded from below, it may lead to vacuum decay via tunnelling. However, this scenario is physically unacceptable, for no late time equilibrium state of the scalar field can be reached then. The effective potential will become negative and huge, creating a large backreaction to eventually push the universe towards a runaway disaster in the anti-de Sitter phase.
@@
We have depicted the behaviour of the non-perturbative effective potential for massless fields, 71, in 8, with $\lambda>16g^{2}$, 73. Let us compare them with the pure scalar field theory results, 2, 3. For a scalar with zero or positive rest mass squared, 3 are rather similar to the tree level potential. 8 on the other hand, shows behaviour qualitatively similar to that of 2, which corresponds to negative rest mass square for the scalar field in a pure $\phi^{4}$-theory. 9 shows the feature of the non-perturbative effective potential with $\lambda=16g^{2}$. Note that for this equality, the $H\to 0$ limit of the effective potential is just the tree level quartic potential, which is very different from that of 9. The above discussion shows that the non-perturbative quantum effects can bring in novel features in spacetimes like the de Sitter. We also recall that for very high energy phenomenon occurring at very small length scales one usually ignores gravity and spacetime curvature. Then the comparison of our non-perturbative results with that of the $H\to 0$ limit also gives us one example of how quantum effects can be very different at short and large scales. Note also that the spontaneous symmetry breaking feature of the flat spacetime (or short scale, sub-Hubble or UV) limit of the effective potential might lead to the formation of non-trivial vacuum structures like the topological defects. For 7, since we essentially have the breaking of the $v\to-v$ reflection symmetry, the corresponding defect will be the domain walls [81]. Also, as we have mentioned above, increasing the Yukawa coupling may make the potential unbounded from below for $\lambda<16g^{2}$. Hence until this limit is reached, increasing the $g$ value with a fixed $\lambda$ would deepen the minima of 7. Thus we may expect that such defect formation would be aided by the Yukawa interaction within this limit. However we also note that, since the full large scale potential washes away the symmetry breaking pattern, it also rules out any such defect formation at that scale. The bound $\lambda\gtrsim 16g^{2}$ found above thus ensures the stability of the vacuum at large scale. This once again strongly emphasises the energy scale dependence of the physical phenomenon we are considering. In other words, formation of topological defects for 71 is possible only if the field is extremely high energetic, and hence confined to a radius of size much small compared to the Hubble horizon, but not otherwise.
@@
Figure 10: (Left) Variation of the effective potential 71, with respect to the Yukawa coupling $g$ for $v/H=2.0$. (Right) $\langle v^{2}\rangle_{\rm stoch.}$ vs. $g$ plot (in the unit of $H^{2}$), assuming the background scalar becomes stochastic at late times. $\lambda$ is taken to be $0.6$ in both plots. Since the fermion mass is $M=gv$, the above plots effectively capture the variation with respect to $M$ as well. See main text for discussion.
Let us also emphasise here the effect of the fermion mass ($M_{\rm dyn.eff}=M_{0}+gv$) in this context. Since we have explicitly focussed on the case $M_{0}=0$ here, the fermion mass contribution comes as $g^{2}v^{2}$ in 71. This makes the effective potential and the action ${\cal Z}_{2}$ symmetric under the reflection $v\to-v$, as has been depicted in the figures. Retaining fermion rest mass however, would break this reflection symmetry, 70. This way the fermion mass can have qualitatively new effect on the correlation functions. We further note that the Yukawa term, including the fermion mass contributes to the effective dynamically generated mass of the scalar field, 59. This effectively leads to the negative rest mass squared like behaviour of 2. Note that for the rest mass squared positive, the behaviour of the effective potential is rather smooth, 3. We have depicted the variation of the effective potential with respect to the Yukawa coupling in the first of 10. Decrease of $V_{\rm eff.loc}$ with increasing $g$-value is consistent with the bound 73. We will further come to this issue in the following Section.
Finally, a possible caveat to our above analysis, as we have mentioned it earlier as well, is the estimation of the value of the secular logarithm numerically, rather than employing any formal resummation procedure (cf., discussion below 54). We wish to attempt now assigning a formal, non-perturbative value to this $\ln a$, should we agree to consider an independent equation for resummation, other than the ones that naturally arise in the 2PI formalism, and thereby simultaneously using two different methods. We adopt the renormalisation group inspired autonomous equation for resummation of de Sitter secular logarithms proposed a couple of years ago in [47]. This method was applied recently in [65], in the context of the $\phi^{4}$-Yukawa theory as follows. At late times, and under the slow roll condition, the vacuum expectation value of the coincident two point scalar propagator satisfies at the leading order,
$\frac{d\langle\phi^{2}\rangle}{d{\cal N}}=\frac{H^{2}}{4\pi^{2}}-\frac{2}{3H^{2}}\left(\frac{\lambda}{6}\langle\phi^{4}\rangle+g\langle\bar{\psi}\psi\phi\rangle\right),$ (74)
where ${\cal N}=\ln a=Ht$ is the number of de Sitter e-foldings. We will take $\langle\phi^{4}\rangle=3\langle\phi^{2}\rangle^{2}$ at the leading order in the Hartree approximation. Note that there is no explicit appearance of dynamical mass in 74. This is because in this method the same is generated via $\langle\phi^{2}\rangle$. Note next that the secular logarithm term in 69 arises from the vacuum expectation value of the local part of the Yukawa term, $\bar{\psi}\psi\phi$, 41. 74 thus suggests that this secular logarithm can be related to $\langle\phi^{2}\rangle$. It also suggests that for an equilibrium state to exist ($d\langle\phi^{2}\rangle/d{\cal N}=0$), the $\ln a$ term cannot be arbitrarily large. Following [47], we next construct the autonomous and non-perturbative equation for $\langle\phi^{2}\rangle$, by eliminating the $\ln a$ terms. It turns out that in the equilibrium, the consistent, non-perturbative value associated with the secular logarithm in question is given by [65],
$\ln a\to\frac{2\sqrt{3}\pi}{\sqrt{\lambda+6g^{2}}},$ (75)
which we may substitute in 69. In other words, we are basically replacing the $\ln a$ term by the non-perturbative value of $\langle\phi^{2}\rangle$, i.e. the scalar vacuum bubble. It is clear that this will bring in no qualitative difference to the result we found earlier by the crude estimation of $\ln a$. However note that in this analysis, we have ignored the background scalar field $v$, and hence the conclusion holds good as long as $\bar{v}=v/H\ll 1$. We may next include $v$ as a rest mass term, $m_{0}^{2}\langle\phi^{2}\rangle$ ($m_{0}^{2}=\lambda v^{2}/2$), on the right hand side of 74. Constructing the autonomous equation as earlier, it is easy to see that we have an additional $\lambda\bar{v}^{2}/2$ term added with $\lambda+6g^{2}$ in the denominator of 75. Thus for $\bar{v}\gg 1$, the $\ln a$ term in 69 will be suppressed compared to the $f_{\rm fin}$ term. Once again this is qualitatively similar to our earlier result, as in the curly bracket of 70, the term multiplied with $0.7$ (originating from $\ln a$ following our numerical estimation) is subleading compared to the $f_{\rm fin}$ term.
Another rigorous approach to resum $\langle\bar{\psi}\psi\phi\rangle$ would be trying to construct the 3PI effective action generating the three point and three particle irreducible effective action, generating a Schwinger-Dyson-like equation for $\langle\bar{\psi}\psi\phi\rangle$. Even though seemingly an extremely formidable task in de Sitter, it seems to be worth attempting.